in the reaction 4NH3(g)+5O2(g)-> 4NO(g)+ 6H2O(l)
a. How many grams of NH3 react with 73.4 g of O2?
b. How many grams of NO are produced from 73.4g of O2?
1
Expert's answer
2020-03-24T14:01:37-0400

"4NH_3(g)+5O_2(g)\to4NO(g)+ 6H_2O(l)"

Moles of "O_2" in "73.4\ g=\frac{73.4}{32}=2.29"

a. "5" moles of "O_2" reacts with "4" moles of "NH_3" .

"2.29" moles of "O_2" will react with "\frac{4}{5}\times 2.29=1.835" moles"=1.835\times 17=31.195\ g"

b. "5" moles of "O_2" produce "4" moles of "NO."

"2.29" moles of "O_2" will produce "\frac{4}{5}\times 2.29=1.835" moles"=1.835\times (14+16)\ g=55.05\ g"